OpenStudy (anonymous):

Solve= -5(-m + 3) = -10m + 30

OpenStudy (anonymous):

-m+3=2m-6 -3m=-9 m=3

OpenStudy (anonymous):

Answer is 3. -5(-m+3)=-10+30 5m-15=-10+30 5m=-10m+45 15=45 A=3